OpenStudy (anonymous):

How do you solve: 2/3= x+7/ 3x

Nnesha (nnesha):

is it \[\frac{ x+7 }{ 3x } \rm ~or~x+\frac{ 7 }{ 3x }\]

OpenStudy (anonymous):

the first one

Nnesha (nnesha):

alright good \[\huge\rm \frac{ 2 }{ 3 }=\frac{ x+7 }{ 3x }\] cross multiply

Nnesha (nnesha):

here is an example \[\huge\rm \frac{ a }{ b }=\frac{ c }{ d}\]\[\rm ad=bc\]

OpenStudy (anonymous):

Do you know how @Amineh0923

OpenStudy (anonymous):

Hint: 2/3 = x+7/3x 2(3x) = 3(x+7)

OpenStudy (anonymous):

Now distribute...

OpenStudy (anonymous):

that's better ;)

OpenStudy (anonymous):

so was I right? its 6x / 3x+7 ? @twistnflip

OpenStudy (anonymous):

eek, try again: 2/3 = x+7/3x 2(3x) = 3(x+7) You are missing one thing try one more time

OpenStudy (anonymous):

6x= 3x+21?

OpenStudy (anonymous):

there you go, awesome. Now do you know how to solve for x?

OpenStudy (anonymous):

I think, I add 6x and 3x because they are like terms which leaves me with 9x = 21, then I divide each side by... oops that's not right :(

OpenStudy (anonymous):

yep you are close though. instead (this is how i was taught) alway cosine terms by subtracting the smallest term from the biggest term. so here subtract 3x from 6x

OpenStudy (anonymous):

combine not cosine

OpenStudy (anonymous):

so.. 6x= 3x + 21 -3x -3x 3x= 21 /3 /3 x=7?

OpenStudy (anonymous):

there you go!! awesome job

OpenStudy (anonymous):

thanks so much! I would fail this on the test if u hadn't cleared that up ;) :P

OpenStudy (anonymous):

no problem anything else? post a new question